The Glass Bottle Company (GBC) manufactures brown glass beverage containers that are sold to breweries. One of the key characteristics of these bottles is their volume. GBC knows that the standard deviation of volume is 0.05 oz. They wish to ensure that the mean volume is not more than 12.10 oz using a sample size of 25 and a level of significance of 0.01. Suppose 25 bottles are measured and the sample mean is 12.15 oz. What is the P-value?

a) 0.563 b) 0.324 c) 0 d)1
Answer #1

please rate, if it helps.. thank you
Answer #2

SOLUTION :


SE = SD of sample / sqrt(n) = 0.05/sqrt(25) = 0.01


t = (sample mean - population mean) / SE = (12.15 - 12.10) / 0.01 = 5


P-value for t = 5 , is 0.000021 = 0  (from p-value calculator) 


Hence, option c : 0 is the ANSWER.

answered by: Tulsiram Garg